The amplifier shows the force F = 60 N. Find the velocity with which the stone with mass = 400 g turns in the circle with radius R = 30 cm.


1
Expert's answer
2021-01-16T17:24:55-0500

Since the stone turns in the circle, the net force acting on it is a centripetal force. According to the second Newton's law, it is equal to:


"F = ma"

where "m = 400g = 0.4kg" is the mass of the stone, and "a" is the centripetal acceleration. "a" is given as follows:


"a = \dfrac{v^2}{R}"

where "v" is the speed of the stone, and "R = 30cm = 0.3m" is the radius of the circle. Subsituting "a" into the expression for "F", and expressing "v", obtain:


"F = m\dfrac{v^2}{R}\\
v = \sqrt{\dfrac{RF}{m}}\\
v = \sqrt{\dfrac{0.3m\cdot 60N}{0.4kg}} \approx 6.7m/s"

Answer. 6.7 m/s.